# Metrics on completely positive maps via noncommutative geometry

### v1

@@ 5.1 Stability
By Lemma 2.7 we have that $F\odot G$ extend to a completely positive map between both $A\otimes_{\rm max}C$ and $B\otimes_{\rm max}D$, and $A\otimes_{\rm min}C$ and $B\otimes_{\rm min}D$. Hence it remains to check that $(\tau_{B}\otimes\tau_{D})((F\otimes G)(1_{A}\otimes 1_{C}))=1$ in both these cases. Indeed,
$(\tau_{B}\otimes\tau_{D})((F\otimes G)(1_{A}\otimes 1_{C}))=\tau_{B}(F(1_{A}))\tau_{D}(G(1_{C}))=1.$
This completes the proof. ∎
@@
Let $A$ be a $C^{*}$-algebra, $B$ be a unital $C^{*}$-algebra, $\delta$ a $C^{*}$-norm on $A\odot B$, $L$ be a seminorm on $A$, and $K$ be a seminorm on $B\otimes_{\delta}A$. If the following conditions are satisfied:
  1. 1.
@@
$\operatorname{mk}_{L}(\varphi_{1},\varphi_{2})=\operatorname{mk}_{K}(\psi\otimes\varphi_{1},\psi\otimes\varphi_{2}),$ (5.1)
@@
$=\sup\{|(\psi\otimes(\varphi_{1}-\varphi_{2}))(1_{n}\otimes x)|\colon x\in A,\ L(x)\leq 1\}$
@@
where the inequality follows because $K(1_{B}\otimes x)\leq 1$ if $L(x)\leq 1$. Next we compute that,
@@
$L((\psi\otimes\operatorname{id}_{A})(X))\leq 1_{B}\otimes L(X)\leq K(X).$
@@
Then the sequence $\{\Delta_{\operatorname{Tr}_{n}\otimes\tau,L_{n}}^{\rm max}\}_{n\in\mathbb{N}}$, where $\Delta_{\operatorname{Tr}_{n}\otimes\tau,L_{n}}^{\rm max}$ is a metric on $\operatorname{TC}_{\operatorname{Tr}_{n}\otimes\tau}(M_{n}(\mathbb{C})\otimes A,M_{n}(\mathbb{C})\otimes B)$, is stable.
@@
To set ourselves in a situation where we can use Lemma 5.2 we define for each $n\geq 2$ a seminorm $K_{n}\colon M_{n}(\mathbb{C})\otimes M_{n}(\mathbb{C})^{\rm op}\otimes A\otimes_{\rm max}B^{\rm op}\to[0,\infty]$ by setting $K_{n}=L_{n}\circ\Sigma_{[23]}$. By conditions 1. and 2. in the hypothesis we get that
  1. 1.
@@
$\operatorname{mk}_{K_{n}}(\varphi,\psi)=\operatorname{mk}_{L_{n}}(\Sigma_{[23]}^{*}\varphi,\Sigma_{[23]}^{*}\psi).$ (5.2)
@@
$=\operatorname{mk}_{L_{n}}(\Sigma_{[23]}^{*}(\omega_{\operatorname{Tr}_{n}}^{\rm max}(\operatorname{id}_{n})\otimes\omega_{\tau}^{\rm max}(F)),\Sigma_{[23]}^{*}(\omega_{\operatorname{Tr}_{n}}^{\rm max}(\operatorname{id}_{n})\otimes\omega_{\tau}^{\rm max}(G)))$ by Equation 4.5,
$=\operatorname{mk}_{K_{n}}(\omega_{\operatorname{Tr}_{n}}^{\rm max}(\operatorname{id}_{n})\otimes\omega_{\tau}^{\rm max}(F),\omega_{\operatorname{Tr}_{n}}^{\rm max}(\operatorname{id}_{n})\otimes\omega_{\tau}^{\rm max}(G))$ by Equation 5.2,
@@
5.2 Chaining
As mentioned in Section 3, we want to talk about chaining for trace channels, and it is not always the case the the composition of two trace channels is again a trace channel. We can however establish that the composition of two trace channels is again a trace channel in the following cases. If $B$ and $C$ are $C^{*}$-algebras having faithful traces $\tau_{B}$ and $\tau_{C}$, respectively, we set
@@ 5.2 Chaining
Let $A$, $B$, and $C$ be $C^{*}$-algebras, where $A$ and $B$ are unital, and $B$ and $C$ have faithful traces $\tau_{B}$ and $\tau_{C}$, respectively. Then composition of linear maps restricts to maps
@@
With this result in mind, we say that a pair $(F,G)$ is composable if
@@
Note that $\operatorname{UCP}(A,B)$ is a subset of $\operatorname{TC}_{\tau_{B}}(A,B)$ only if $\tau_{B}$ is a tracial state. This will only impact the wording of Theorem 5.10, where all traces will be states.
Before we can state the main results of this section, we are going to need some technical results. This first result says how the map $\omega_{\tau}$ interacts with compositions.
@@
$=(\omega_{\tau}(G)\circ(F\odot\operatorname{id}_{C^{\rm op}}))(a\otimes c^{\rm op}).$
If $L$ is a seminorm on $A\otimes_{\rm max}B^{\rm op}$, then we get a seminorm $L^{\rm op}$ on $B\otimes_{\rm max}A^{\rm op}$ by setting $L^{\rm op}=L\circ\Sigma^{\rm op}$. The following result says that this is the “correct” seminorm to consider if we want the distance between the adjoint of two channels to be the same as the distance between the original channels.
@@
By Equation 4.8 and continuity of $\omega_{\tau_{B}}(F)$ and $\omega_{\tau_{A}}(F^{\sharp})$ in the respective maximal norms, we get that
@@
By Section 2.2: if $L_{A}$ and $L_{B}$ are seminorms on $A$ and $B$ respectively, then we have that
$(L_{A\otimes_{\rm max}B^{\rm op}})^{\rm op}=L_{B\otimes_{\rm max}A^{\rm op}}.$ (5.3)
The following Theorem and Corollary 5.12, are the main results of this section. Especially Corollary 5.12 is needed to establish that chaining holds for the class of unital completely positive maps between group $C^{*}$-algebras that arise from positive definite functions, cf. Theorem 6.9.
@@
$\leq\operatorname{mk}_{L^{\rm op}}(\omega_{\tau_{A}}^{\rm max}(F_{1}^{\sharp}),(\omega_{\tau_{A}}^{\rm max}(F_{2}^{\sharp}))$ $\text{since }L^{\rm op}\circ(G_{2}^{\sharp}\otimes\operatorname{id}_{A^{\rm op}})\leq M^{\rm op}$
@@
If the traces $\tau_{A}$, $\tau_{B}$, and $\tau_{C}$ are amenable, then the above result has a version where the maximal tensor product is interchanged with the minimal tensor product instead. The proof of this is completely analogous to the one given above, so we omit the details here.
@@
Let $A$, $B$, and $C$ be unital $C^{*}$-algebras with faithful traces $\tau_{A}$, $\tau_{B}$, and $\tau_{C}$ respectively, and with seminorms $L_{A}$, $L_{B}$, and $L_{C}$ respectively. Suppose that
  1. 1.
@@
Note that all the pairs $(F_{1},G_{1})$, $(F_{1},G_{2})$, and $(F_{2},G_{2})$ are composable, hence we are in the setting of Theorem 5.10. By Remark 5.9 we simply need to show that the inequalities $L_{B}\circ F_{1}\leq L_{A}$ and $L_{B}\circ G_{2}^{\sharp}\leq L_{C}$ imply that
@@
6 Applications
6.1 The Kasparov exterior product of spectral triples
Recall that we defined unital spectral triples in Section 2.2. In this section we will show that using the Kasparov exterior product from $KK$-theory, we can from any spectral triple generate a sequence of seminorms on the matrix amplifications which naturally satisfy attractive stability properties as discussed in Section 3. This will be made precise below, see Theorem 6.2. We proceed to closely follow the exposition given in [Kaa24]. Let $(A,H_{A},\partial_{A})$ be a spectral triple, and suppose $\mathcal{B}\subseteq B(H_{B})$ be a unital *-subalgebra. We may then form the essentially self-adjoint operators
@@ 6.1 The Kasparov exterior product of spectral triples
$\begin{split}\partial_{A}\widehat{\otimes}1_{\mathcal{B}}&\colon{\rm Dom}(\partial_{A})\otimes H_{B}\to H_{A}\otimes H_{B}\\ 1_{\mathcal{B}}\widehat{\otimes}\partial_{A}&\colon H_{B}\otimes{\rm Dom}(\partial_{A})\to H_{B}\otimes H_{A}\end{split}$ (6.1)
While $\partial_{A}\widehat{\otimes}\operatorname{id}_{H_{B}}$ and $\operatorname{id}_{H_{B}}\widehat{\otimes}\partial_{B}$ do not generally have compact resolvents as self-adjoint unbounded operators on $H_{A}\otimes H_{B}$ and $H_{B}\otimes H_{A}$, respectively, we can still use them to induce metrics much in the same way as in Equation 2.2. More precisely, by identifying elements of the algebraic tensor products with operators on the Hilbert space tensor products in the obvious way, we could set
@@
$\begin{split}L_{\partial_{A}\widehat{\otimes}1_{\mathcal{B}}}=L_{\partial_{A}}\otimes 1_{B}\quad\text{and}\quad L_{1_{\mathcal{B}}\widehat{\otimes}\partial_{A}}=1_{B}\otimes L_{\partial_{A}}\end{split}$ (6.3)
@@
We proceed to discuss the exterior Kasparov product of unital spectral triples. Let therefore $(A,H_{A},\partial_{A})$ and $(B,H_{B},\partial_{B})$ be unital spectral triples. We may form the exterior Kasparov product $(A\otimes_{\rm min}B,H,\partial_{A}\times\partial_{B})$, where the Hilbert space $H$, and the self-adjoint operator $\partial_{A}\times\partial_{B}$ depend on the parity of the two spectral triples as follows:
@@
Let $(A,H_{A},\partial_{A})$ and $(B,H_{B},\partial_{B})$ be spectral triples. Then for any $x\in A\otimes_{\rm min}B$ we have that
@@
for every $x\in A\otimes_{\rm min}B^{\rm op}$ satisfying $L_{\partial_{A}\times\partial_{B}}(x)\leq 1$. To prove this we need to consider the parity of the three spectral triples $(A,H_{A},\partial_{A})$, $(B^{\rm op},H_{B},\partial_{B})$, and $(M_{n}(\mathbb{C}),H_{n},\partial_{n})$. Denote the resulting external product by
$((M_{n}(\mathbb{C})\otimes M_{n}(\mathbb{C})^{\rm op})\otimes(A\otimes_{\rm min}B^{\rm op}),H,\partial).$
The proof is a straightforward but tedious exercise in writing out Kasparov products. There are eight cases to consider, and in every case one verifies that $L_{\partial}(1_{n}\otimes 1_{n}\otimes x)=L_{\partial_{A}\times\partial_{B}}(x)$ for all $x\in A\otimes_{\rm min}B^{\rm op}$. We write out what $H$ and $\partial$ is in the following exemplary case.
Even, even, and even: In this case $H=(H_{n}\otimes H_{n})\otimes(H_{A}\otimes H_{B})$, and
$\partial=(\partial_{n}\times\partial_{n})\otimes 1+(\gamma_{n}\otimes\gamma_{n})\otimes(\partial_{A}\times\partial_{B}).$
@@
$=\|(\gamma_{n}\otimes\gamma_{n})\otimes[\partial_{A}\times\partial_{B},x]\|$
$=\|(\gamma_{n}\otimes\gamma_{n})\|\|[\partial_{A}\times\partial_{B},x]\|$
@@
$=L_{\partial_{A}\times\partial_{B}}(x).$
Even though $\partial$ will look more complicated in some of the other cases, it is still relatively straightforward to conclude that $L_{\partial}(1_{n}\otimes 1_{n}^{\rm op}\otimes x)=L_{\partial_{A}\times\partial_{B}}(x)$ for all $x\in A\otimes_{\rm min}B^{\rm op}$. This completes the proof. ∎
@@
Note that one could use two different spectral triples for $M_{n}(\mathbb{C})$ and for $M_{n}(\mathbb{C})^{\rm op}$ in Theorem 6.2. We have chosen to not do this for the ease of notation.
6.2 Group $C^{*}$-algebras and length functions
By specializing to countable discrete groups, we will in this section construct seminorms on twisted group $C^{*}$-algebras from which we can construct stable metrics on the set of trace channels, see Example 6.10. Moreover, there is a distinguished class of unital completely positive maps on twisted group $C^{*}$-algebras coming from the positive definite functions on the group, and we show that these maps behave well with respect to the chaining property for the same metrics, see Theorem 6.9.
Suppose now $G$ is a discrete group and that $\sigma\in Z^{2}(G,\mathbb{T})$ is a $2$-cocycle. The *-algebra $C_{c}(G,\sigma)$ consists of the complex functions on $G$ with finite support, equipped with convolution and involution defined by
$\delta_{g}(x)=\begin{cases}1&x=g,\\ 0&x\neq g.\end{cases}$
for $f\in C_{c}(G,\sigma)$, $\xi\in\ell^{2}(G)$ and $x\in G$. This defines a faithful *-representation of the opposite algebra $C^{*}_{r}(G,\sigma)^{\rm op}$
Definition 6.4 ([Con89]).
  1. 1.
    $l(s)=0$ if $s=1_{G}$,
  2. 2.
    $l(s^{-1})=l(s)$ for all $s\in G$, and
  3. 3.
    $l(st)\leq l(s)+l(t)$ for all $s,t\in G$.
Suppose moreover that $G$ is countable and the length function $l$ is proper, that is, for every $r\geq 0$, the set $\ell^{-1}([0,r])$ is a finite set. Such a length function gives rise to an essentially self-adjoint unbounded operator $\partial_{l}$ on $\ell^{2}(G)$ given by the linear extension of
$\begin{split}\partial_{l}\colon C_{c}(G)&\to\ell^{2}(G)\\ \delta_{s}&\mapsto l(s)\delta_{s}\end{split}$
As $\partial_{l}$ is essentially self-adjoint, we may take its self-adjoint closure, which we will also denote by $\partial_{l}$ to ease notation. We thus obtain a spectral triple $(C^{*}_{r}(G,\sigma),\ell^{2}(G),\partial_{l})$, where $C^{*}_{r}(G,\sigma)$ is faithfully represented on $\ell^{2}(G)$ with the left regular representation of $C_{c}(G,\sigma)$. We furthermore obtain a spectral triple $(C^{*}_{r}(G,\sigma)^{\rm op},\ell^{2}(G),\partial_{l})$, where $C^{*}_{r}(G,\sigma)^{\rm op}$ is faithfully represented on $\ell^{2}(G)$ with the right regular representation of $C_{c}(G,\sigma)$. Note that a straightforward calculation such as in [Rie02, p. 614] shows that for any $f$ in $C_{c}(G,\sigma)$, the commutator $\partial_{l},\lambda^{\sigma}(f)]$ belongs to the $\sigma$-twisted crossed product $\ell^{\infty}(G)\rtimes_{{\rm red},\sigma}G$, being here identified with the $C^{*}$-subalgebra of $B(\ell^{2}(G))$ generated by $\ell^{\infty}(G)$ (acting as multiplication operators) and $\{\lambda^{\sigma}_{g}:g\in G\}$.
Let $P(G)$ denote the positive definite functions on $G$, and set
[…]
is a completely positive map such that $\|M_{\phi}\|=\phi(1_{G})$ (resp. a unital completely positive map), see e.g. [BO08, Theorem 2.5.11] (when $\sigma=1$) or [BC16, Lemma 2.6].

### v2

@@ 6.1. Stability
By lemma 2.7 we have that $F\odot G$ extends to a completely positive map between both $A\otimes_{\rm max}C$ and $B\otimes_{\rm max}D$, and $A\otimes_{\rm min}C$ and $B\otimes_{\rm min}D$. Hence, it remains only to check that $(\tau_{B}\otimes\tau_{D})((F\otimes G)(1_{A}\otimes 1_{C}))=1$ in both these cases. Indeed, we have
$(\tau_{B}\otimes\tau_{D})((F\otimes G)(1_{A}\otimes 1_{C}))=\tau_{B}(F(1_{A}))\tau_{D}(G(1_{C}))=1.\qed$
@@
Let $A$ be a $C^{*}$-algebra, $B$ be a unital $C^{*}$-algebra, $\delta$ be a $C^{*}$-norm on $B\odot A$, $L$ be a seminorm on $A$, and $K$ be a seminorm on $B\otimes_{\delta}A$. Assume the following conditions are satisfied:
  1. (1)
@@
(6.3) $\operatorname{mk}_{L}(\varphi_{1},\varphi_{2})=\operatorname{mk}_{K}(\psi\otimes\varphi_{1},\psi\otimes\varphi_{2}),$
@@
$=\sup\{|(\psi\otimes(\varphi_{1}-\varphi_{2}))(1_{B}\otimes x)|\colon x\in A,\ L(x)\leq 1\}$
@@
where the inequality follows by assumption (2). Next we compute that
@@
$L((\psi\otimes\operatorname{id}_{A})(X))\leq(1_{B}\otimes L)(X)\leq K(X).$
@@
Then the sequence $\{\Delta_{\operatorname{Tr}_{n}\otimes\tau,L_{n}}^{\rm max}\}_{n\in\mathbb{N}}$ is stable, where $\Delta_{\operatorname{Tr}_{n}\otimes\tau,L_{n}}^{\rm max}$ is an extended metric on $\operatorname{TC}_{\operatorname{Tr}_{n}\otimes\tau}(M_{n}(\mathbb{C})\otimes A,M_{n}(\mathbb{C})\otimes B)$.
@@
To set ourselves in a situation where we can use lemma 6.3 we define for each $n\geq 2$ a seminorm $K_{n}\colon M_{n}(\mathbb{C})\otimes M_{n}(\mathbb{C})^{\rm op}\otimes A\otimes_{\rm max}B^{\rm op}\to[0,\infty]$ by setting $K_{n}=L_{n}\circ\Sigma_{[23]}$. By conditions (1) and (2) in the statement of the theorem we get that
  1. (1)
@@
(6.4) $\operatorname{mk}_{K_{n}}(\varphi,\psi)=\operatorname{mk}_{L_{n}}(\Sigma_{[23]}^{*}\varphi,\Sigma_{[23]}^{*}\psi).$
@@
$=\operatorname{mk}_{L_{n}}(\Sigma_{[23]}^{*}(\omega_{\operatorname{Tr}_{n}}^{\rm max}(\operatorname{id}_{n})\otimes\omega_{\tau}^{\rm max}(F)),\Sigma_{[23]}^{*}(\omega_{\operatorname{Tr}_{n}}^{\rm max}(\operatorname{id}_{n})\otimes\omega_{\tau}^{\rm max}(G)))$ by (5.5),
$=\operatorname{mk}_{K_{n}}(\omega_{\operatorname{Tr}_{n}}^{\rm max}(\operatorname{id}_{n})\otimes\omega_{\tau}^{\rm max}(F),\omega_{\operatorname{Tr}_{n}}^{\rm max}(\operatorname{id}_{n})\otimes\omega_{\tau}^{\rm max}(G))$ by (6.4),
@@
6.2. Chaining
As mentioned at the beginning of this section, we want to talk about chaining for trace channels, but it is not always the case the composition of two trace channels is again a trace channel. We can however establish that the composition of two trace channels is again a trace channel in the following cases. If $B$ and $C$ are $C^{*}$-algebras having faithful traces $\tau_{B}$ and $\tau_{C}$, respectively, we set
@@ 6.2. Chaining
Let $A$, $B$, and $C$ be $C^{*}$-algebras, where $A$ and $B$ are unital, and $B$ and $C$ have faithful traces $\tau_{B}$ and $\tau_{C}$, respectively. Then the composition of linear maps restricts to maps
@@
With this result in mind, we say that a pair $(F,G)$ is $\operatorname{TC}$-composable if
@@
Note that $\operatorname{UCP}(A,B)$ is a subset of $\operatorname{TC}_{\tau_{B}}(A,B)$ if (and only if) $\tau_{B}$ is a tracial state. Also, $\operatorname{QC}(B,C,\tau_{B},\tau_{C})$ is a subset of $\operatorname{TC}_{\tau_{C}}(B,C)$ whenever $\tau_{B}$ is a tracial state. This will only impact the wording of theorem 6.11 and corollary 6.13, where all traces will be states.
Before we can state the main results of this section, we are going to need some technical results. This first result tells us how the map $\omega_{\tau}$ interacts with compositions.
@@
$=(\omega_{\tau}(G)\circ(F\odot\operatorname{id}_{C^{\rm op}}))(a\otimes c^{\rm op}).\qed$
If $L$ is a seminorm on $A\otimes_{\rm max}B^{\rm op}$, then we get a seminorm $L^{\rm op}$ on $B\otimes_{\rm max}A^{\rm op}$ by setting $L^{\rm op}=L\circ\Sigma^{\rm op}$. The following result says that this is the “correct” seminorm to consider if we want the distance between the adjoints of two channels to be the same as the distance between the original channels.
@@
By (5.9) and continuity of $\omega_{\tau_{B}}(F)$ and $\omega_{\tau_{A}}(F^{\sharp})$ in the respective maximal norms, we get that
@@
It follows from section 2.2 that if $L_{A}$ and $L_{B}$ are seminorms on $A$ and $B$, respectively, then we have
(6.5) $(L_{A\otimes_{\rm max}B^{\rm op}})^{\rm op}=L_{B\otimes_{\rm max}A^{\rm op}}.$
The following theorem and corollary 6.13 are the main results of this section. Especially corollary 6.13 is needed to establish that chaining holds for the class of unital completely positive maps between group $C^{*}$-algebras that arise from positive definite functions, cf. theorem 7.7.
@@
$\leq\operatorname{mk}_{L^{\rm op}}(\omega_{\tau_{A}}^{\rm max}(F_{1}^{\sharp}),\omega_{\tau_{A}}^{\rm max}(F_{2}^{\sharp}))$ $\text{since }L^{\rm op}\circ(G_{2}^{\sharp}\otimes\operatorname{id}_{A^{\rm op}})\leq M^{\rm op}$
@@
If the tracial states $\tau_{A}$, $\tau_{B}$, and $\tau_{C}$ are amenable, then the above result has a version where the maximal tensor product is interchanged with the minimal tensor product instead. The proof of this is completely analogous to the one given above, so we omit the details here.
@@
Let $A$, $B$, and $C$ be unital $C^{*}$-algebras with faithful tracial states $\tau_{A}$, $\tau_{B}$, and $\tau_{C}$, respectively, and with seminorms $L_{A}$, $L_{B}$, and $L_{C}$, respectively. Suppose that
  1. (1)
@@
Note that all the pairs $(F_{1},G_{1})$, $(F_{1},G_{2})$, and $(F_{2},G_{2})$ are $\operatorname{TC}$-composable, hence we are in the setting of theorem 6.11. By remark 6.10 we simply need to show that the inequalities $L_{B}\circ F_{1}\leq L_{A}$ and $L_{B}\circ G_{2}^{\sharp}\leq L_{C}$ imply that
@@
7. Applications
7.1. The Kasparov external product of spectral triples
Recall that we defined unital spectral triples in section 2.2. In this section we will show that using the Kasparov external product from $KK$-theory, we can from any spectral triple generate a sequence of seminorms on the matrix amplifications which naturally satisfy attractive stability properties, as discussed in section 6.1. This will be made precise below, see theorem 7.2. We proceed to closely follow the exposition given in [Kaa24]. Let $(A,H_{A},\partial_{A})$ be a unital spectral triple, and suppose that $\mathcal{B}\subseteq B(H_{B})$ is a unital *-subalgebra. We may then form the essentially self-adjoint operators
@@ 7.1. The Kasparov external product of spectral triples
(7.1) $\begin{split}\partial_{A}\widehat{\otimes}1_{\mathcal{B}}&\colon{\rm Dom}(\partial_{A})\otimes H_{B}\to H_{A}\otimes H_{B}\\ 1_{\mathcal{B}}\widehat{\otimes}\partial_{A}&\colon H_{B}\otimes{\rm Dom}(\partial_{A})\to H_{B}\otimes H_{A}\end{split}$
While $\partial_{A}\widehat{\otimes}1_{\mathcal{B}}$ and $1_{\mathcal{B}}\widehat{\otimes}\partial_{B}$ do not generally have compact resolvents as self-adjoint unbounded operators on $H_{A}\otimes H_{B}$ and $H_{B}\otimes H_{A}$, respectively, we can still use them to induce metrics much in the same way as in (2.2). More precisely, by identifying elements of the algebraic tensor products with operators on the Hilbert space tensor products in the obvious way, we could set
@@
(7.3) $\begin{split}L_{\partial_{A}\widehat{\otimes}1_{\mathcal{B}}}=L_{\partial_{A}}\otimes 1_{B}\quad\text{and}\quad L_{1_{\mathcal{B}}\widehat{\otimes}\partial_{A}}=1_{B}\otimes L_{\partial_{A}}\end{split}$
@@
We proceed to discuss the external Kasparov product of unital spectral triples. Let therefore $(A,H_{A},\partial_{A})$ and $(B,H_{B},\partial_{B})$ be unital spectral triples. We may form the external Kasparov product $(A\otimes_{\rm min}B,H,\partial_{A}\times\partial_{B})$, where the Hilbert space $H$, and the self-adjoint operator $\partial_{A}\times\partial_{B}$ depend on the parity of the two spectral triples as follows:
@@
Let $(A,H_{A},\partial_{A})$ and $(B,H_{B},\partial_{B})$ be unital spectral triples. Then for any $x\in A\otimes_{\rm min}B$ we have that
@@
for every $x\in A\otimes_{\rm min}B^{\rm op}$ satisfying $L_{\partial_{A}\times\partial_{B}}(x)\leq 1$. We will in fact prove the stronger statement that
(7.4) $L_{\partial}(1_{n}\otimes 1_{n}^{\rm op}\otimes x)=L_{\partial_{A}\times\partial_{B}}(x)$
for all $x\in A\otimes_{\rm min}B^{\rm op}$. To prove this we only need to consider the parity of the spectral triple for $M_{n}(\mathbb{C})\otimes M_{n}(\mathbb{C})^{\rm op}$, and the parity of the spectral triple for $A\otimes_{\rm min}B^{\rm op}$. In fact, since we are considering the same spectral triple for $M_{n}(\mathbb{C})$ and $M_{n}(\mathbb{C})^{\rm op}$, the resulting external product of these two spectral triples will always be even. Hence, we only need to consider the parity of the spectral triple for $A\otimes_{\rm min}B^{\rm op}$. In both cases we see that
$\partial=(\partial_{n}\times\partial_{n})\otimes 1+\gamma_{M_{n}(\mathbb{C})\otimes M_{n}(\mathbb{C})^{\rm op}}\otimes(\partial_{A}\times\partial_{B}).$
@@
$=\|[(\partial_{n}\times\partial_{n})\otimes 1+\gamma_{M_{n}(\mathbb{C})\otimes M_{n}(\mathbb{C})^{\rm op}}\otimes(\partial_{A}\times\partial_{B}),1_{n}\otimes 1_{n}^{\rm op}\otimes x]\|$
$=\|[\gamma_{M_{n}(\mathbb{C})\otimes M_{n}(\mathbb{C})^{\rm op}}\otimes(\partial_{A}\times\partial_{B}),1_{n}\otimes 1_{n}^{\rm op}\otimes x]\|$
$=\|\gamma_{M_{n}(\mathbb{C})\otimes M_{n}(\mathbb{C})^{\rm op}}\otimes[\partial_{A}\times\partial_{B},x]\|$
$=\|\gamma_{M_{n}(\mathbb{C})\otimes M_{n}(\mathbb{C})^{\rm op}}\|\|[\partial_{A}\times\partial_{B},x]\|$
@@
$=L_{\partial_{A}\times\partial_{B}}(x).\qed$
@@
Note that one could use two different spectral triples for $M_{n}(\mathbb{C})$ and for $M_{n}(\mathbb{C})^{\rm op}$ in theorem 7.2. If one chooses to do this in a way such that these spectral triples have different parities then one would have to check that equation (7.4) still holds in this case.
7.2. Amenable groups and twisted group $C^{*}$-algebras
By specializing to countable amenable discrete groups, we will in this section construct a stable sequence of extended metrics on the set of trace channels on a twisted group $C^{*}$-algebra, see example 7.8. Moreover, we will show that the class of unital completely positive maps on twisted group $C^{*}$-algebras coming from the normalized positive definite functions on the group behave well with respect to the chaining property for the same metrics, see theorem 7.7. We use the notation introduced in section 2.3.